OpenStudy (anonymous):

Solve. 9x2 – 21x + 15 = 0.

hero (hero):

9x2 – 21x + 15 = 0 3(3x^2 - 7x + 5) = 0 3x^2 - 7x + 5 = 0 Use the quad formula and let a = 3, b = -7, c = 5 \[x = \frac{-b \pm \sqrt{b^2 - 4ac}}{2a}\]

OpenStudy (anonymous):

I still don't understand how to do it? I sqaure 21 and get 441 - 540. What do I do now?